mhchen:

Do you guys agree with this valid math statement: 0.2 = 0.19999999...

mhchen:

\[0.1999...=0.1+\frac{9}{100}(1+\frac{1}{10}+\frac{1}{10^{2}}+...)=0.1+\frac{9}{100}\sum_{i=0}^{\infty}(\frac{1}{10})^{i}\] and \[\sum_{n=1}^{\infty}(r)^{n}=\frac{1}{1-r}\] then \[0.1999...=0.1+\frac{9}{100}\frac{1}{1-\frac{1}{10}}=0.1+0.1=0.2\] lol

Hero:

Easier to just do this: \( \begin{array}{l} 10x\ =1.9999...\\ 100x\ =\ 19.9999...\\ \\ 100x\ -\ 10x\ =\ 18\\ 90x\ =\ 18\\ x=\dfrac{18}{90}\\ =\dfrac{1}{5}\\ =0.2 \end{array}\)

mhchen:

Wow that's so much simpler

Hero:

indeed

mhchen:

Imagine how annoying that'd be if instead of writing 3 + 2 = 5 I'd start writing 2.99999... + 1.9999... = 4.99999.. and give my answer to teachers ahaha